OpenStudy (anonymous):

find the slope of the tangent to the graph of y = sqrt(x+2) at x= -1 PLEASE SHOW ALL STEPS THANK YOU! :)

jimthompson5910 (jim_thompson5910):

\[\Large y(x) = \sqrt{x+2}\] \[\Large y^{\prime}(x) = \frac{d}{dx}(\sqrt{x+2})\] \[\Large y^{\prime}(x) = \frac{1}{2\sqrt{x+2}}\] \[\Large y^{\prime}(-1) = \frac{1}{2\sqrt{-1+2}}\] \[\Large y^{\prime}(-1) = \frac{1}{2\sqrt{1}}\] \[\Large y^{\prime}(-1) = \frac{1}{2(1)}\] \[\Large y^{\prime}(-1) = \frac{1}{2}\] So the slope of the tangent line at x = -1 is \[\Large \frac{1}{2}\]